HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025  Not logged in ELOG logo
Message ID: 82     Entry time: Wed Apr 2 15:07:52 2025
Author: Michael, Alex, Maeve 
Category: Accelerator 
Subject: Still not a good beam in the ESR 
The (power supply of) SIS kicker magnet has been repaired and the beam has been restored from about 14:20.
But, the Schottky spectrum in the ESR looked pretty bad, so the measurement can not be restarted. 
The accelerator experts are currently looking into this problem.
ELOG V3.1.5-fc6679b